Flexible budget for selling and administrative expenses for a service company Digital Solutions Inc. uses flexible...

  1. Flexible budget for selling and administrative expenses for a service company

    Digital Solutions Inc. uses flexible An accounting device used to plan and control resources of operational departments and divisions.budgets that are based on the following data:

    Sales commissions 9% of sales
    Advertising expense 16% of sales
    Miscellaneous administrative expense $11,500 per month plus 3% of sales
    Office salaries expense $48,000 per month
    Customer support expenses $16,000 per month plus 33% of sales
    Research and development expense $79,000 per month

    This information has been collected in the Microsoft Excel Online file. Open the spreadsheet, perform the required analysis, and input your answers in the question below.

    Open spreadsheet

    Prepare a flexible selling and administrative expenses budget for October for sales volumes of $400,000, $900,000, and $1,000,000. (Use Exhibit 5 as a model.) Enter all amounts as positive numbers.

    Digital Solutions Inc.
    Flexible Selling and Administrative Expenses Budget
    For the Month Ending October 31
    Total sales $400,000 $900,000 $1,000,000
    Variable cost:
    • Cash
    • Office salaries expense
    • Sales commissions
    • Supplies expense
    • Wages expense
    $ $ $
    • Advertising expense
    • Cash
    • Office salaries expense
    • Supplies expense
    • Wages expense
            
    • Miscellaneous administrative expense
    • Office salaries expense
    • Prepaid insurance
    • Rent expense
    • Wages expense
            
    • Customer support expenses
    • Rent expense
    • Research and development expense
    • Supplies expense
    • Wages expense
            
    Total variable cost $ $ $
    Fixed cost:
    • Insurance expense
    • Miscellaneous administrative expense
    • Rent expense
    • Supplies expense
    • Wages expense
    $ $ $
    • Insurance expense
    • Office salaries expense
    • Prepaid insurance
    • Rent expense
    • Wages expense
            
    • Customer support expenses
    • Insurance expense
    • Rent expense
    • Supplies expense
    • Wages expense
            
    • Insurance expense
    • Rent expense
    • Research and development expense
    • Supplies expense
    • Wages expense
            
    Total fixed cost $ $ $
    Total selling and administrative expenses $ $ $

    Prepare a flexible selling and administrative expenses budget for October for sales volumes of $400,000, $900,000, and $1,000,000. (Use Exhibit 5 as a model.) Enter all amounts as positive numbers.

    Digital Solutions Inc.
    Flexible Selling and Administrative Expenses Budget
    For the Month Ending October 31
    Total sales $400,000 $900,000 $1,000,000
    Variable cost:
    $ $ $
            
            
            
    Total variable cost $ $ $
    Fixed cost:
    $ $ $
            
            
            
    Total fixed cost $ $ $
    Total selling and administrative expenses $ $ $

    Prepare a flexible selling and administrative expenses budget for October for sales volumes of $400,000, $900,000, and $1,000,000. (Use Exhibit 5 as a model.)

    Digital Solutions Inc.
    Flexible Selling and Administrative Expenses Budget
    For the Month Ending October 31
    Total sales $400,000 $900,000 $1,000,000
    Variable cost:
    Sales commissions $36,000 $81,000 $90,000
    Advertising expense    64,000    144,000    160,000
    Miscellaneous administrative expense    12,000    27,000    30,000
    Customer support expenses    132,000    297,000    330,000
    Total variable cost $244,000 $549,000 $610,000
    Fixed cost:
    Miscellaneous administrative expense $11,500 $11,500 $11,500
    Office salaries expense    48,000    48,000    48,000
    Customer support expenses    16,000    16,000    16,000
    Research and development expense    79,000    79,000    79,000
    Total fixed cost $154,500 $154,500 $154,500
    Total selling and administrative expenses $398,500 $703,500 $764,500
Answer #1

Answer

  • Requirement

Digital Solutions Inc.

Flexible Selling and Administrative Expenses Budget

For the Month Ending October 31

Total sales

$400,000

$900,000

$1,000,000

Variable cost:

Sales commissions

$36,000

$81,000

$90,000

Advertising expense

$64,000

$144,000

$160,000

Miscellaneous administrative expense

$12,000

$27,000

$30,000

Customer support expenses

$132,000

$297,000

$330,000

Total variable cost

$244,000

$549,000

$610,000

Fixed cost:

Miscellaneous administrative expense

$11,500

$11,500

$11,500

Office salaries expense

$48,000

$48,000

$48,000

Customer support expenses

$16,000

$16,000

$16,000

Research and development expense

$79,000

$79,000

$79,000

Total fixed cost

$154,500

$154,500

$154,500

Total selling and administrative expenses

$398,500

$703,500

$764,500

--Working for variable cost

Total sales

400000

900000

1000000

Variable cost:

Sales commissions

=400000*9%

=900000*9%

=1000000*9%

Advertising expense

=400000*16%

=900000*16%

=1000000*16%

Miscellaneous administrative expense

=400000*3%

=900000*3%

=1000000*3%

Customer support expenses

=400000*33%

=900000*33%

=1000000*33%